# DriveGPT: Scaling Autoregressive Behavior Models for Driving

### v2

@@ 5.1.3 Closed-Loop Driving
We demonstrate the effectiveness of DriveGPT as a real-time motion planner deployed in a closed-loop setting. The model takes input features from an industry-level perception system that outputs agent states and map information.
In Fig. 10, we present a challenging example in dense urban traffic where there are two double-parked vehicles blocking the path forward along with other oncoming vehicles.  DriveGPT generates smooth and safe trajectories, bypassing the blocking vehicles and moving back to the original lane afterward. More examples are presented in Appendix C.1.
Figure 10: Example of DriveGPT running as a closed-loop planner at real time in dense urban traffic. It first produces a forward trajectory (top), updates it to bypass double-parked vehicles (middle), and drives back to the original lane at the end (bottom).
@@ 5.2 External Evaluation: Motion Prediction
HDGT [8] 0.7676 1.1077 0.1325 0.3709
MTR [22] 0.6050 1.2207 0.1351 0.4216
HPTR [29] 0.5565 1.1393 0.1434 0.3904
Wayformer Ensemble [17] 0.5454 1.1280 0.1228 0.4335
MotionLM Ensemble [21] 0.5509 1.1199 0.1058 0.4507
@@
Table 4: On the WOMD test set,  DriveGPT achieves better results in geometric metrics without any ensembling, and improved performance after pretraining. Results are averaged over three agent types (vehicle, pedestrian, cyclist) and three prediction horizons (3s, 5s, 8s). The best metric is highlighted in bold and the second best is underlined.
@@ 5.2.1 Open-Source Encoder
For our external evaluation, we use the open-source MTR [22] encoder. This encoder is similar to the one described in Sec. 3.2. We make this change to improve the reproducibility of our results and take advantage of MTR’s open-source dataloading code for WOMD. We use the same autoregressive decoder as described in Sec. 3.3.
@@ 5.2.3 Results
We present two variants of our method to validate its effectiveness on the motion prediction task, including DriveGPT-WOMD that is trained on WOMD, and DriveGPT-Finetune that is pretrained on our 120M internal research dataset and finetuned on WOMD. For baselines, we use a set of representative state-of-the-art models reported in [21].
We report results on the WOMD test set in Table 4. The results demonstrate that our method outperforms existing state-of-the-art non-ensemble models in terms of geometric metrics. Compared to Wayformer [17] and MotionLM [21] that use ensembles of up to 8 replicas, our model achieves the best minADE and minFDE metrics and the second-best miss rate metric without any ensembling.
While we prioritize geometric metrics that emphasize the recall of predicted trajectory samples (i.e. not missing the critical trajectory), our model shows lower soft mAP scores due to suboptimal probability estimates. These estimates suffer from accumulated noises over time, as they are computed by compounding the action probabilities over a long sequence (as described in Eq. (1)). Consequently, we notice that the gap in soft mAP grows as the prediction horizon increases. This reveals a limitation of using an autoregressive decoder for accurate probability estimation, as also noted in the LLM literature [10, 5]. We defer improving the probability estimates of autoregressive models for behavior modeling as future work.
@@ 6 Conclusion
We introduced DriveGPT, an LLM-style autoregressive behavior model, to better understand the effects of model parameters and dataset size for autonomous driving. We systematically explored model performance as a function of both dataset size and model parameters, uncovering LLM-like scaling laws for data and compute scaling, along with limited model scaling properties. We showed the quantitative and qualitative benefits of scaling for planning in real-world driving scenarios. Additionally, we demonstrated our method on a public motion prediction benchmark, where DriveGPT outperformed a state-of-the-art baseline by a large margin and achieved improved performance through pretraining on a large-scale dataset.
@@ A.1 Internal Evaluation: AV Planning
We train our models on the internal research dataset for 1 epoch, as customary in the LLM scaling literature [11, 6].
Our models follow the implementations described in [17, 21], and are trained using a batch size of 2048 and a standard Adam optimizer adopted in [11]. We follow the optimal learning rate schedule discovered in [6], which applies a cosine decay with a cycle length equivalent to the total number of training steps. All models were trained on 16 NVIDIA H100 GPUs.
@@ A.2 External Evaluation: Motion Prediction
We train our models on the WOMD data following the same setup in [22]. More specifically, we use a batch size of 80 and an AdamW optimizer with a learning rate of 0.0001. The models are trained for 30 epochs, where the learning rate is decayed by a factor of 0.5 every 2 epochs, starting from epoch 20.
Appendix B Scaling Detail
B.1 Scaling Configurations
In our scaling experiments, we vary the model size through the hidden dimension $d_{\text{model}}$ in both encoder and decoder transformers, as summarized in Table. 6. For each model size, we run multiple experiments at different maximum learning rates, and present the optimal value in Table. 6.
B.2 Ablation Study on Attention Heads
@@
Model Size Hidden Dimension ($d_{\text{model}}$) Max LR
1.5M 64 0.0050
4M 128 0.0020
8M 192 0.0014
12M 256 0.0010
26M 384 0.0010
94M 768 0.0007
163M 1024 0.0004
358M 1536 0.0002
629M 2048 0.0001
1.4B 3072 0.0001
Table 6: We vary the model size over three orders of magnitude through hidden dimensions. For each model size, we report the optimal learning rate, which decreases as the model size increases, matching the observations in the LLM scaling literature [11, 6].
@@
B.3 Ablation Study on Decoder Scaling
@@
B.4 Ablation Study on Encoder Scaling
@@
B.5 Scaling From a Perplexity View
@@
Appendix C Additional Qualitative Examples
@@
C.1 Additional Closed-Loop Driving Examples
In the supplementary video [^2] https://youtu.be/-hLi44PfY8g, we present additional examples of using DriveGPT as a real-time motion planner in a closed loop setting. The video covers representative challenging scenarios in driving, including a) unprotected left turn, b) double parked vehicle, c) construction zone, d) blow-through cyclist, and e) lane change in heavy traffic.
C.2 Failure Cases on WOMD
Appendix D Additional WOMD Validation Results
D.1 WOMD Data Scaling in Pretraining
D.2 WOMD Model Scaling

### v3

@@ 5.1.3 Closed-Loop Driving
We demonstrate the effectiveness of DriveGPT as a real-time motion planner deployed in a closed-loop setting. The model takes input features from an industry-level perception system that outputs agent states and map information. We used the 8M DriveGPT model, trained on the full dataset, to drive the car, achieving a latency of under 50ms on a single onboard GPU.
In Fig. 10, we present a challenging example in dense urban traffic where there are two double-parked vehicles blocking the path forward along with other oncoming vehicles.  DriveGPT generates smooth and safe trajectories, bypassing the blocking vehicles and moving back to the original lane afterward. More examples are presented in the supplementary video at https://youtu.be/-hLi44PfY8g, where DriveGPT alone is responsible for driving in real time.
Figure 10: Example of DriveGPT running as a closed-loop planner in real time in dense urban traffic. It first produces a forward trajectory (top), updates it to bypass double-parked vehicles (middle), and drives back to the original lane at the end (bottom).
@@ 5.2 External Evaluation: Motion Prediction
MTR (Shi et al., 2022) 0.6050 1.2207 0.1351 0.4216
HDGT (Jia et al., 2023) 0.7676 1.1077 0.1325 0.3709
HPTR (Zhang et al., 2023) 0.5565 1.1393 0.1434 0.3968
ControlMTR (Sun et al., 2024) 0.5897 1.1916 0.1282 0.4572
MTR++ (Shi et al., 2024) 0.5906 1.1939 0.1298 0.4414
Wayformer (Nayakanti et al., 2023) 0.5454 1.1280 0.1228 0.4335
MotionLM (Seff et al., 2023) 0.5509 1.1199 0.1058 0.4507
@@
Table 5: On the WOMD test set,  DriveGPT achieves better results in geometric metrics without any ensembling, and improved performance after pretraining. Results are averaged over three agent types (vehicle, pedestrian, cyclist) and three prediction horizons (3s, 5s, 8s). The best metric is highlighted in bold and the second best is underlined. $\dagger$ denotes ensemble.
@@ 5.2.1 Open-Source Encoder
For our external evaluation, we use the open-source MTR (Shi et al., 2022) encoder. This encoder is similar to the one described in Sec. 3.2. We make this change to improve the reproducibility of our results and take advantage of MTR’s open-source dataloading code for WOMD. We use the same autoregressive decoder as described in Sec. 3.3.
@@ 5.2.3 Results
We present two variants of our method to validate its effectiveness on the motion prediction task, including DriveGPT-WOMD that is trained on WOMD, and DriveGPT-Finetune that is pretrained on our 120M internal research dataset and finetuned on WOMD. For baselines, we use a set of representative state-of-the-art models.
We report results on the WOMD test set [^2] Metrics are sourced from (Zhang et al., 2023) and WOMD leaderboard. in Table 5. The results demonstrate that our method outperforms existing state-of-the-art non-ensemble models in terms of geometric metrics. Compared to Wayformer (Nayakanti et al., 2023) and MotionLM (Seff et al., 2023) that use ensembles of up to 8 replicas, our model achieves the best minADE and minFDE metrics and the second-best miss rate metric without any ensembling.
While we prioritize geometric metrics that emphasize the recall of predicted trajectory samples (i.e. not missing the critical trajectory), our model shows lower soft mAP scores due to suboptimal probability estimates. These estimates suffer from accumulated noises over time, as they are computed by compounding the action probabilities over a long sequence (as described in Eq. (1)), resulting in less accurate sample probabilities and reduced soft mAP, which relies on accurate probability assignments across predicted samples. Consequently, we notice that the gap in soft mAP grows as the prediction horizon increases. This reveals a limitation of using an autoregressive decoder for accurate probability estimation, as also noted in the LLM literature (Jiang et al., 2021; Geng et al., 2024). We defer improving the probability estimates of autoregressive models for behavior modeling as future work. One potential direction is to train an additional probability prediction head for each sample, which could enhance probability estimates and lead to improved soft mAP scores.
@@ 6 Conclusion
We introduced DriveGPT, an LLM-style autoregressive behavior model, to better understand the effects of model parameters and dataset size for autonomous driving. We systematically examined model performance as a function of both dataset size and model capacity, revealing LLM-like scaling laws for data and compute, as well as diminishing returns with increased model size. We showed the quantitative and qualitative benefits of scaling for planning in real-world driving scenarios. Additionally, we demonstrated our method on a public motion prediction benchmark, where DriveGPT outperformed state-of-the-art baselines and achieved improved performance through pretraining on a large-scale dataset.
Acknowledgments
We thank, in alphabetical order, Abbas Shikari, Akshay Rangesh, Allan Lazarovici, Andy Zhuang, Ann Yan, Anton Spiridonov, Burkay Donderici, David Fang, Dennis Park, Florent Bekerman, Greg Meyer, Hans Zhang, Haytham Mohamed, Joe Liu, Jonathan Sorg, Junyuan Shi, Ling Li, Mao Ye, Matthew Creme, Peng Tang, Poonam Suryanarayan, Rizwan Chaudhry, Rosalie Zhu, Siva Karthik Mustikovela, Srinidhi Myadaboyina, Tanmay Agarwal, Xinlei Pan, Yi Hao, Yunlong He, Zaiwei Zhang, Zisu Dong, and many others from Cruise data and infra teams, for their contributions to the paper.
Impact Statements
This paper aims to advance research in behavior modeling for autonomous vehicles. Since AVs interact with the public, our work has the potential to enhance their behavior around other road users, ultimately contributing to safer roads.
@@ A.1 Internal Evaluation: AV Planning
We train our models on the internal research dataset for 1 epoch, as customary in the LLM scaling literature (Kaplan et al., 2020; Hoffmann et al., 2022).
Our models follow the implementations described in (Nayakanti et al., 2023; Seff et al., 2023), and are trained using a batch size of 2048 and a standard Adam optimizer adopted in (Kaplan et al., 2020). We follow the optimal learning rate schedule discovered in (Hoffmann et al., 2022), which applies a cosine decay with a cycle length equivalent to the total number of training steps. All models were trained on 16 NVIDIA H100 GPUs.
@@ A.2 External Evaluation: Motion Prediction
We train our models on the WOMD data following the same setup in (Shi et al., 2022). More specifically, we use a batch size of 80 and an AdamW optimizer with a learning rate of 0.0001. The models are trained for 30 epochs, where the learning rate is decayed by a factor of 0.5 every 2 epochs, starting from epoch 20.
Appendix B Additional Ablation Studies
B.1 Ablation Study on Attention Heads
@@
B.2 Ablation Study on Decoder Scaling
@@
B.3 Ablation Study on Encoder Scaling
@@
Appendix C Scaling From a Perplexity View
@@
Appendix D Additional Qualitative Examples
D.1 Qualitative Comparison of Large Models
In Fig. 16, we present two qualitative examples demonstrating that a larger DriveGPT model (163M parameters) exhibits better collision avoidance capabilities, compared to its smaller 94M-parameter variant.
Figure 16: A larger DriveGPT model at 163M parameters generates trajectories (red) that maintain a safer distance from other traffic agents and road debris compared to a smaller model at 94M parameters (pink).
@@ D.1 Qualitative Comparison of Large Models
D.2 Additional Closed-Loop Driving Examples
In the supplementary video at https://youtu.be/-hLi44PfY8g, we present additional examples of deploying DriveGPT as a real-time motion planner in a closed-loop setting. The video covers representative challenging scenarios in driving, including a) unprotected left turn, b) double parked vehicle, c) construction zone, d) blow-through cyclist, and e) lane change in heavy traffic.
D.3 Failure Cases on WOMD
Appendix E Additional WOMD Validation Results
E.1 WOMD Data Scaling in Pretraining
E.2 WOMD Model Scaling